OpenStudy (anonymous):

What is the 19th term of the arithmetic sequence 14, 22, 30, … ? 158 150 174 166

OpenStudy (anonymous):

d=8

OpenStudy (anonymous):

8*18+14

OpenStudy (anonymous):

Tn = 14+(n-1)*8 T(19) = 14+(19-1)*8

OpenStudy (anonymous):

158?

OpenStudy (anonymous):

+ 8 38, 46, 54, 62, etc

OpenStudy (anonymous):

158 is correct.